OpenStudy (anonymous):

What is the solution to the system? y = 2x + 7 y = x - 1

geerky42 (geerky42):

We know that y = 2x + 7 and y = x - 1, so 2x + 7 = x - 1 Now find x.

OpenStudy (anonymous):

I dont know how to fined x

geerky42 (geerky42):

Just isolate x. Do you know how to?

OpenStudy (anonymous):

yes

geerky42 (geerky42):

Do it.

OpenStudy (anonymous):

I got 8 = x and sorry my computer is not wanting to work

OpenStudy (anonymous):

does 8=x? @geerky42

OpenStudy (nubeer):

hmm its not 8 .. it's -8

OpenStudy (anonymous):

oh thank you

OpenStudy (nubeer):

now just put this value of x in the equation you will get the y

OpenStudy (anonymous):

so the answer is (-8,-9)

OpenStudy (anonymous):

Thank you @geerky42 and @nubeer!!!